OpenStudy (anonymous):

What is the solution of the following system? -3x - 2y = -12 9x + 6y = -9

OpenStudy (austinl):

Is there a particular method we are supposed to use?

OpenStudy (anonymous):

It does not state a method :(

OpenStudy (anonymous):

I sense a problem with this system. After multiplying the top equation by 3 and adding the two together, it kinda falls apart.

OpenStudy (anonymous):

Solving both equations for y to get slope-intercept form will also point out this particular feature.

OpenStudy (anonymous):

My available answer choices are; no solutions, infinitely many solutions, (2,1) or (-2,-1)

OpenStudy (austinl):

Ok, in that case, elimination is probably the simplest and easiest route for this one. You would multiply the top equation by 3 to get \(-9x-6y=-36\) You would add that together with the bottom equation. \(\cancel{-9x}-\cancel{6y}=-36\) \(\cancel{9x}+\cancel{6y}=-9\) \(0\ne-45\) No solution.

OpenStudy (austinl):

In all honesty, this problem was mostly busy work....

OpenStudy (anonymous):

Busy work + knowledge of a concept: It's important to know what 'no solution' means in this context, i.e. the two lines are parallel (have identical slope) and therefore do not intersect.

OpenStudy (anonymous):

Thank you all so much for your help!!!

OpenStudy (austinl):

True, if you solve them both out for \(y\) you should be able to see this. \(-3x - 2y = -12\\ -2y=3x-12\\ y=\color{blue}{-\dfrac{3}{2}x}+6\) \(9x + 6y = -9\\ 6y=-9x-9\\ y=\color{blue}{-\dfrac{3}{2}x}-9\)